### v2

Abstract
Network reconfiguration can significantly increase the hosting capacity (HC) for distributed generation (DG) in radially operated systems, thereby reducing the need for costly infrastructure upgrades. However, when the objective is DG maximization, jointly optimizing topology and power dispatch remains computationally challenging. Existing approaches often rely on relaxations or approximations, yet we provide counterexamples showing that interior point methods, linearized DistFlow and second-order cone relaxations all yield erroneous results. To overcome this, we propose a solution framework based on the exact DistFlow equations, formulated as a bilinear program and solved using spatial branch-and-bound (SBB). Numerical studies on standard benchmarks and a 533-bus real-world system demonstrate that our proposed method reliably performs reconfiguration and dispatch within time frames compatible with real-time operation.
1 Introduction
Distribution networks are generally built in a looped fashion, but operated with one point normally-open in each loop resulting in radial networks. The set of switches to leave open (hence the network configuration) offers a degree-of-freedom that is normally not exploited in the industry. In this paper, we study the joint optimization of network reconfiguration and power control of distributed generation (DG), and we call this the reconfiguration optimal power flow (ROPF) problem, generalizing optimal power flow (OPF). While enforcing network configuration and operational constraints (e.g., voltage and line current limits and DG restrictions), the ROPF problem traditionally seeks to minimize objectives including total line loss, fuel cost or voltage deviation (e.g., [3, 15, 33, 1, 24, 2]). However, in this paper we consider the maximization of total DG output in ROPF as an emerging topic for renewable energy integration (e.g., [5, 10, 27, 13]). The DG maximization objective has a crucial effect on the tractability of the ROPF problem and the usefulness of the resulting dispatch solution. This is the first main discovery of this paper.
Diverse solution methods for the ROPF problem have been explored. For instance, heuristics and rule-based approaches have been considered due to their ease for online computation (e.g., [3, 1, 12]). To more explicitly search for the optimal solution, the ROPF problem is formulated and solved as an optimization problem, akin to OPF’s solution approach. A variety of methods have been proposed to solve the ROPF optimization problem. Meta-heuristics algorithms (e.g., [32, 35, 10]) are commonly considered due to their flexibility and ease of use. However, lack of optimality guarantee and unpredictable runtime are major drawbacks of these methods. Thus, more direct optimization methods are explored. For example, using DC power flow equations or linearized DistFlow equations to approximate the power system physics, [23, 9, 6, 26] investigate mixed integer linear programming (MILP) models for ROPF. Alternatively, to directly include the AC power flow equations, mixed integer conic programming (MICP) ROPF models are proposed (e.g., [15, 8, 33, 17]). Theoretical support and numerical demonstrations have been reported (e.g., [8, 20, 15, 33, 19, 13]). However, it is not well recognized that the objective of the ROPF problem plays an important role in the suceess of the solution approaches. In particular, except for [13] all aformentioned previous works consider “classical” objectives such as loss minimization, convex quadratic fuel cost minimization, conservation voltage reduction, etc. For the “non-classical” objective of DG maximization in this paper, it turns out that popular methods such as the MICP apporach and interior point algorithm may lead to undesirable control dispatch violating grid constraints such as voltage and current limits. We support our claim with a repeatable case study involving a three-bus example with all computation details presented. To the best of our knowledge, the adverse effects of these pitfalls have not been well recognized in the literature.
The difficulty with DG maximizing ROPF necessitates a solution method with acceptable accuracy and efficiency. With the only nonlinearity in the model being polynomials (due to the power flow equations), the ROPF problem can be interpreted as a bilinear programming problem using specialized solver (e.g., Gurobi). Similar approaches have been studied in [18, 5], where the ROPF problem is treated as a more general mixed integer nonlinear programming (MINLP) problem and solved using the spatial branch-and-bound (SBB) algorithm. In terms of accuracy, the SBB algorithm is sufficient since it guarantees optimality regardless of the ROPF objective type. However, since SBB is fundamentally an exhaustive search it is important to investigate its computational efficiency for the DG maximizing ROPF problem. The rather long runtime reported in [18, 5] does not imply efficiency especially for real-time applications. The efficiency of SBB is significantly affected by the choice of problem formulation and algorithm implementation. A crucial factor in problem formulation is the model of power system physics described by the power flow equations. Typical forms of the power flow equations include, for example, AC power flow equations with voltages in polar coordinates (e.g., [18]) and voltages in rectangular coordinates (e.g., [5]), branch flow equations [8], DistFlow equations [3], modified AC power flow equations (e.g., [7, 16]), etc. On the other hand, different SBB implementations (e.g., Gurobi, BARON, Knitro) behave differently because they make different assumptions about the problem structure. Through our case studies, we identify the combination of DistFlow equations and Gurobi (version 9 or later, with recently added capability to handle bilinear programs) as a promising choice of solution method for the DG maximizing ROPF problem. For instance, for a 533-bus real example in Sweden, ROPF can be performed within 10 minutes when allowing up to four switch status changes. To the best of our knowledge, successful demonstration of SBB to solve ROPF models with the exact DistFlow equations is not well-known, even though the DistFlow equations are commonly relaxed and used in the MICP approach (e.g., [8]). We emphasize that the appropriate choice of formulation and SBB solver is important. We will provide counterexamples to illustrate the poor computational performance of some improper choices. We will also demonstrate that the DistFlow equations play a vital role in the efficient computation of the ROPF problem. This is the second main discovery of this paper.
The main novelties and contributions of this paper are listed as follows:
  • We raise the awareness that the DG maximizing objective fundamentally changes the tractability of the ROPF problem and illustrate the pitfalls of possible overextension of popular solution methods such as MICP relaxation, interior point algorithms and linearization approximation.
  • We point out the importance of the choice of problem formulation and optimization solver for the DG maximizing ROPF problem, by illustrating a promising choice and other improper choices.
The rest of this paper is organized as follows: Section 2 introduces the notation, defines the ROPF optimization model and describes the proposed solution approach. Section 3 illustrates and evaluates the performance of the proposed ROPF approach. Section 4 justifies the proposed method by showing pitfalls encountered by common alternatives for DG maximizing ROPF. Section 5 discusses a practical renewable energy hosting capacity maximization problem for a 533-bus distribution network in Sweden using the proposed method.
2 Problem Description
2.1 Problem Overview
We consider balanced power distribution systems with line shunt elements ignored so that the system can be described by the DistFlow equations [3]. The lines are equipped with switches which can be closed or open. The network is built with loop(s) but is operated in a radial topology. Power exchange is possible at the slack bus. In addition, load and DG are present in the system. The loads are assumed to be known. The DG sources are interfaced with inverters where both active and reactive power outputs are considered adjustable. By jointly adjusting network configuration and inverter outputs, the ROPF problem considered in this paper seeks to maximize the total distributed active power generation in the system with the following constraints: (a) DistFlow equations, (b) network radiality and allowable switch status changes, (c) inverter power output bounds, apparent power capacity, power factor and (d) grid constraints (voltage and line current limits).
We comment that our adopted DistFlow model, while restricted by the three-phase balance assumption, is well accepted as evidenced by numerous influential results built upon the same assumption (e.g., [3, 15, 33]). Also, it is common to find representative distribution system models described exactly by the DistFlow equations (e.g., MATPOWER [36], REDS repository [1], EU examples [22]).
2.2 Notation and Assumptions
[…]
  • For any bus $n\in\bar{\mathcal{N}}$, the voltage is $v_{n}$. We define $\nu_{n}:=|v_{n}|^{2}$ as the squared voltage for bus $n$. $\nu_{n}$ is a state decision variable in the ROPF problem.
  • For any bus $n\in\mathcal{N}$, the active and reactive power loads are $p^{L}_{n}$ and $q^{L}_{n}$. Their values are assumed to be known.
  • For any bus $n\in\mathcal{N}$, the active and reactive power generation are $p^{G}_{n}$ and $q^{G}_{n}$. They are control decision variables of the ROPF problem.
  • In a radial network, each bus $n\in\mathcal{N}$ is associated with exactly one line, denoted by line $n$, with reference direction pointing away from bus 0. The current and power flow follow the same reference direction of the line.
  • $\pi_{n}\in\bar{\mathcal{N}}$ is the “parent” bus of $n\in\mathcal{N}$ (so that line $n$ goes from $\pi_{n}$ to $n$).
  • For any $n\in\bar{\mathcal{N}}$, $c(n)\subseteq\mathcal{N}$ is the set of buses with parent being $n$. That is, $c(n)=\{k\in\mathcal{N}\mid\pi_{k}=n\}$.
  • For any bus $n\in\mathcal{N}$, $P_{n}$ is the active power flow from bus $\pi_{n}$ to bus $n$ measured at bus $\pi_{n}$. $Q_{n}$ is defined similarly for reactive power flow. $P_{n}$ and $Q_{n}$ are state decision variables of the ROPF problem.
  • Line shunt elements are ignored. For any bus $\in\mathcal{N}, n\ell_{n}$ is the squared magnitude current from bus $\pi_{n}$ to bus $n$ (also the value from $n$ to $\pi_{n}$). $\ell_{n}$ is a state decision variable in the ROPF problem.
  • For a line $\{n,k\}$ in a reconfigurable network, the series resistance and reactance are $r_{nk}$ and $x_{nk}$ respectively. Also, the square line current upper limit is denoted by $\overline{\ell}_{nk}$. Note that $r_{nk}=r_{kn}$, $x_{nk}=x_{kn}$ and $\overline{\ell}_{nk}=\overline{\ell}_{kn}$.
  • For a reconfigurable network, $\pi_{nk}\in\{0,1\}$ with $(n,k)\in\bar{\mathcal{N}}\times\bar{\mathcal{N}}$ such that $\pi_{nk}=1$ if and only if bus $n$ is the parent of bus $k$ (i.e., $\pi_{k}=n$). $\pi_{nk}$ is a ROPF control decision variable. Also, we define the auxiliary decision variable $\alpha_{nk}=\pi_{nk}+\pi_{kn}$ for $(n,k)\in\bar{\mathcal{N}}\times\bar{\mathcal{N}}$ to encode the switch status of the line between bus $n$ and bus $k$.
2.3 ROPF Problem Constraints Derivation
2.3.1 DistFlow Equations for Fixed Network Configuration
With fixed network configuration, the physics of distribution systems is described by the DistFlow equations [3]:
$p^{G}_{n}-p^{L}_{n}$ $=-P_{n}+\sum\limits_{k\in c(n)}P_{k}+r_{n}\ell_{n},$ $n\in\mathcal{N}$ (1a)
$q^{G}_{n}-q^{L}_{n}$ $=-Q_{n}+\sum\limits_{k\in c(n)}Q_{k}+x_{n}\ell_{n},$ $n\in\mathcal{N}$ (1b)
$\nu_{\pi_{n}}-\nu_{n}$ $=2r_{n}P_{n}+2x_{n}Q_{n}-(r_{n}^{2}+x_{n}^{2})\ell_{n},$ $n\in\mathcal{N}$ (1c)
$\nu_{\pi_{n}}\ell_{n}$ $=P_{n}^{2}+Q_{n}^{2},$ $n\in\mathcal{N}$ (1d)
where $r_{n}$ and $x_{n}$ are the series resistance and series reactance of line $n$, respectively. The set $c(n)$ is defined in Section $2.$
2.3.2 Network Configuration Modelling
We impose bounds on the switch status decision variables $\alpha_{nk}$ defined in Section 2.2 to indicate always closed and always open switches:
$\underline{\alpha}_{nk}\leq\alpha_{nk}\leq\overline{\alpha}_{nk},\qquad(n,k)\in\bar{\mathcal{N}}\times\bar{\mathcal{N}}$ (2)
A network with $N+1$ buses is radial if and only if it is connected and contains exactly NN lines. Radiality can be imposed by the (necessary and sufficient) single-commodity flow conditions. Using the parent decision variables $\pi_{nk}$ defined in Section 2.2 and auxiliary continuous “flow” variables $f_{ij}$ for $(i,j)\in\bar{\mathcal{N}}\times\bar{\mathcal{N}}$, these conditions can be described as [28, 29]
$\sum\limits_{k\in\bar{\mathcal{N}}}f_{nk}-\sum\limits_{k\in\bar{\mathcal{N}}}f_{kn}=-1,$ $n\in\mathcal{N}$ (3a)
$0\leq f_{nk}\leq N\pi_{nk},$ $(n,k)\in\bar{\mathcal{N}}\times\bar{\mathcal{N}}$ (3b)
$\sum\limits_{(n,k)\in\bar{\mathcal{N}}\times\bar{\mathcal{N}}}\pi_{nk}=N.$ (3c)
We may wish to restrict the number of switch status changes denoted by $K$. Let parameters $\alpha^{0}_{nk}\in\{0,1\}$ for $(n,k)\in\bar{\mathcal{N}}\times\bar{\mathcal{N}}$ denote the initial network configuration such that $\alpha_{nk}=\alpha_{kn}=1$ if switch $\{n,k\}$ is initially closed and $\alpha_{nk}=\alpha_{kn}=0$ if $\{n,k\}$ is initially open. Then, the desired constraint is
$\sum\limits_{(n,k)\in\bar{\mathcal{N}}\times\bar{\mathcal{N}}}\!\!\Big(\alpha^{0}_{nk}(1-\alpha_{nk})+\alpha_{nk}(1-\alpha^{0}_{nk})\Big)\leq 2K.$ (4)
Note that when a switch is closed another switch must be opened simultaneously to keep the radial network configuration. Hence, $K$ is typically an even number.
2.3.3 DistFlow Equations for Configurable Network
When the network configuration is parameterized by decision variable $\pi$ (also $\alpha$), certain configuration related terms in (1) should be generalized. For instance, the children set $c(n)$ in (1a) and (1b) is not determined before the problem is solved. Thus, the following changes are needed in (1a) and (1b)
$\sum\limits_{k\in c(n)}P_{k}\rightarrow\sum\limits_{k\in\mathcal{N}}\pi_{nk}P_{k},\;\;\sum\limits_{k\in c(n)}Q_{k}\rightarrow\sum\limits_{k\in\mathcal{N}}\pi_{nk}Q_{k},\;\;n\in\mathcal{N}$
to describe the sum of power flows leaving bus $n\in\mathcal{N}$. Similarly, while line $n$ always points to bus $n$, the parent $\pi_{n}$ is not fixed before optimization. The line parameters and parent bus voltage in (1a) to (1d) become
$r_{n}\rightarrow\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}r_{kn},\;\;x_{n}\rightarrow\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}x_{kn},\;\;\nu_{\pi_{n}}\rightarrow\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}\nu_{k},$
where $r_{kn}$ and $x_{kn}$ are the series resistance and reactance of line $\{n,k\}$ as defined in Section $2.$ Hence, the DistFlow equations for configurable networks become
$p^{G}_{n}\!-\!p^{L}_{n}=\!-P_{n}\!\!+\!\!\sum\limits_{k\in\mathcal{N}}\pi_{nk}P_{k}\!+\!\!\Big(\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}r_{kn}\Big)\ell_{n},\;n\!\in\!\mathcal{N}$ (5a)
$q^{G}_{n}\!-\!q^{L}_{n}=\!-Q_{n}\!\!+\!\!\sum\limits_{k\in\mathcal{N}}\pi_{nk}Q_{k}\!+\!\!\Big(\sum\limits_{k\in\bar{\mathcal{N}}}\!\pi_{kn}x_{kn}\Big)\ell_{n},n\!\in\!\mathcal{N}$ (5b)
$\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}\nu_{k}-\nu_{n}=2(\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}r_{kn})P_{n}+2(\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}x_{kn})Q_{n}$
$\quad-\Big(\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}(r_{kn}^{2}+x_{kn}^{2})\Big)\ell_{n},\qquad n\in\mathcal{N},$ (5c)
$\Big(\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}\nu_{k}\Big)\ell_{n}=P_{n}^{2}+Q_{n}^{2},\qquad n\in\mathcal{N}.$ (5d)
Contrary to the fixed configuration setting in (1), the line variables $P_{n}$, $Q_{n}$ and $\ell_{n}$ in (5) do not correspond to a predefined line in the network. These line variables can be associated with any line incident to bus $n$. The actual association is determined by the values of the $\pi$ variables at optimality. We note that the formulation in (5) is different from standard reconfiguration-type formulation in the literature, where big-M type status changes are directly imposed on (5c).
In (5a), the bilinear term $\pi_{nk}P_{k}$ with 0-1 binary variable $\pi_{nk}$ and continuous variable $P_{k}$ can be equivalently replaced by an auxiliary variable with additional linear constraints as follows. Suppose the continuous variable $P_{k}$ is bounded such that $\underline{P}_{k}\leq P_{k}\leq\overline{P}_{k}$. With auxiliary continuous variable $z$ satisfying
$\underline{P}_{k}\pi_{nk}\leq z\leq\overline{P}_{k}\pi_{nk},$ (6a)
$P_{k}-\overline{P}_{k}(1-\pi_{nk})\leq z\leq P_{k}-\underline{P}_{k}(1-\pi_{nk}),$ (6b)
the bilinear term $\pi_{nk}P_{k}$ can be replaced by $z$. The claim can be verified by separately considering the cases when $\pi_{nk}=0$ and $\pi_{nk}=1$. Other binary-continuous bilinear terms $\pi_{kn}\ell_{n}$, $\pi_{nk}Q_{k}$ and $\pi_{kn}\nu_{k}$ in (5) can be replaced by auxiliary variables in a similar fashion. However, in (5d) the terms $(\pi_{kn}\nu_{k})\ell_{n}$ remain bilinear even after the replacement of $\pi_{kn}\nu_{k}$. This complicates the solution process of the ROPF problem.
2.3.4 Inverter Operational Constraints
The inverter operational constraints include (a) generation bounds, (b) apparent power rating, and (c) minimum power factor. These constraints are
$\underline{p}^{G}_{n}\leq p^{G}_{n}\leq\overline{p}^{G}_{n},\quad\underline{q}^{G}_{n}\leq q^{G}_{n}\leq\overline{q}^{G}_{n},$ $n\in\mathcal{N}$ (7a)
$(p^{G}_{n})^{2}+(q^{G}_{n})^{2}\leq(C_{n})^{2},$ $n\in\mathcal{N}$ (7b)
$|q^{G}_{n}|\leq\tan(\cos^{-1}(\text{pf}_{n}))p^{G}_{n},$ $n\in\mathcal{N}$ (7c)
where $\underline{p}^{G}_{n}$, $\overline{p}^{G}_{n}$, $\underline{q}^{G}_{n}$, $\overline{q}^{G}_{n}$, $C_{n}$, $\text{pf}_{n}$ are given parameters for active and reactive power generation bounds, inverter apparent power rating and minimum power factor for bus $n$, respectively.
2.3.5 Grid Constraints
For operational safety, voltage limits are imposed:
$\underline{\nu}_{n}\leq\nu_{n}\leq\overline{\nu}_{n},\qquad n\in\bar{\mathcal{N}}$ (8)
where $\underline{\nu}_{n}$ and $\overline{\nu}_{n}$ are given. Common values are $\sqrt{\underline{\nu}_{n}}=0.95$ pu, $\sqrt{\overline{\nu}_{n}}=1.05$ pu. Also, line current limit constraints are
$0\leq\ell_{n}\leq\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}\overline{\ell}_{kn},\qquad n\in\mathcal{N}$ (9)
where $\overline{\ell}_{kn}$ is the given square current upper limit for line $\{k,n\}$. Typically, $\sqrt{\overline{\ell}_{kn}}$ is up to a few hundred amperes.
The replacement of the bilinear terms $\pi_{nk}P_{k}$ in (5a) and $\pi_{nk}Q_{k}$ in (5b) requires variable bounds on the line power flows $P_{k}$ and $Q_{k}$. If these bounds are not provided, they can be derived as follows:
$-\sqrt{\nu^{U}_{k}{\ell}^{U}_{k}}\leq P_{k}\leq\sqrt{\nu^{U}_{k}{\ell}^{U}_{k}},$ $k\in\mathcal{N}$ (10)
$-\sqrt{\nu^{U}_{k}{\ell}^{U}_{k}}\leq Q_{k}\leq\sqrt{\nu^{U}_{k}{\ell}^{U}_{k}},$ $k\in\mathcal{N}$ (11)
where $\nu^{U}_{k}:=\max\limits_{(n,k)\mid\overline{\alpha}_{nk}=1}\overline{\nu}_{n}$ and ${\ell}^{U}_{k}:=\max\limits_{(n,k)\mid\overline{\alpha}_{nk}=1}\overline{\ell}_{nk}$. The bounds in (10) and (11) follow necessarily from (8) and (9). $P_{k}+jQ_{k}=v_{\pi_{k}}(I_{k})^{*}$ with $v$ and II being the complex voltage and line current. Thus, $|P_{k}|\leq|P_{k}+jQ_{k}|\leq|v_{\pi_{k}}||I_{k}|\leq\sqrt{\nu^{U}_{k}{\ell}^{U}_{k}}$. This establishes (10). The same argument shows (11).
2.4 DG maximizing ROPF Model and the Proposed Solution Approach
Maximizing the total DG active power output, the ROPF problem studied in this paper can be summarized as
[…]
In (12), the decision variables without subscripts denote the vectors or matrices stacked by the corresponding variables for individual buses or lines (e.g., $p^{G}$ is the NN-vector of all $p^{G}_{n}$ for $n\in\mathcal{N}$). $\nu,\ell,p^{G},q^{G},P,Q$ are continuous vector variables, while $\pi$ and $\alpha$ are 0-1 binary matrix variables. Except for (5d), all constraints can be rewritten as linear equalities or inequalities (cf. (6)). In contrast, in (5d) the product of continuous variables $\nu_{k}$ and $\ell_{n}$ cannot be equivalently replaced by any auxiliary variable as in (6) (the transformation requires at least one 0-1 binary variable in the product). Thus, problem (12) is a mixed integer linear program with additional non-convex bilinear equality constraints.
Even though (12) resembles a classical model of the ROPF problem (if the objective is changed to loss minimization), we emphasize that the DG maximization objective in (12) leads to significant ramifications that render most of the typical solution methods ineffective. The details will be provided in Section 4. Instead, we propose to solve (12) using the SBB (i.e., spatial branch-and-bound) algorithm of Gurobi (version 9 or later), and we subsequently call this our “proposed method”. The rationale of our choice is as follows:
  1. (a)
    SBB returns (sub)optimal solutions with certifiable optimality gap.
  2. (b)
    The SBB implementation in Gurobi is specialized in bilinear programming problem, allowing more effective use of the problem structure.
  3. (c)
    Case studies with realistically sized benchmark instances indicate the practical computational performance of the proposed method.
We note that different packages implementing the SBB algorithm (e.g., Gurobi, BARON, Knitro) behave differently for problem (12). The numerical experiment in Section 4.5 indicates that Gurobi is the best performer among our choices.
3 Illustration and Discussion of the Proposed Method
We evaluate the performance of the proposed method by solving instances of the ROPF problem in (12) derived from benchmarks listed in Table 1.
Table 1: Distribution system benchmark data
name # of buses # of tie-switches # of DG source
33bw 33 5 2 [3, 36]
118zh 118 15 10 [36]
136ma 136 21 7 [11, 36]
REDS 29+1 30 1 4 [1]
REDS 83+11 84 13 4 [1, 32]
REDS 135+1 136 21 4 [1, 11]
REDS 201+3 202 15 5 [11, 1]
REDS 873+7 874 27 5 [1]
The topology, line parameters and load data of the benchmarks can be found in the respective sources in Table 1. For each benchmark we add DG sources in the system (location details omitted). The rating of each DG source is 10 MVA for all benchmarks except for REDS 873+7 where the rating is 50 MVA (large enough to prevent the capacity constraint (7b) from being active).
For each bus the active power generation lower and upper limits in (7a) are zero and the DG source apparent power rating, respectively. On the other hand, in (7a) reactive power generation is not sign-restricted, and its absolute value is no greater than the DG rating. For (7c), the minimum power factor is 0.9 for all DG sources. Unless specified otherwise, for all buses the voltage lower and upper limits in (8) are $0.95$ pu and $1.05$ pu, respectively. The current upper limit in (9), on the other hand, varies from line to line. The values are in the order of a few hundred amperes. We also consider a variant of 33bw with line current rating enlarged to 600 A for all lines.
All computation in this and the next section is performed on a PC with an Intel i9-12900K CPU and 64 GB of RAM. The ROPF problem in (12) is solved with Gurobi 10.0.1 (with parameter “non-convex” set to 2) and we use AMPL to implement the optimization models. Gurobi is called with its default parameter values (e.g., target optimality gap 0.01%).
3.1 Reconfiguration with the 33-bus Benchmark
We consider the 33-bus benchmark 33bw described in Table 1. The topology, the locations of the two DG sources and the line current ratings are shown in Fig. 1.
Figure 1: 33-bus distribution system from [3] with five tie-switches (dashed lines). Line width is proportional to current rating labeled next to the line.
We first attempt to solve the reconfiguration problem in (12) maximizing the total DG output with $K=0$ (i.e., OPF with original network topology). The instance is infeasible due to operational constraints. However, if we relax the voltage lower limit from 0.95 pu to 0.9 pu, the instance becomes feasible and the voltages from bus 12 to bus 17 (towards the end of the feeder) are below 0.95 pu. Next, we solve (12) with $K=2$ (i.e., one switch opened and another switch closed) with the usual voltage limits of 0.95 pu and 1.05 pu. The result is shown in Fig. 2, with a total DG output of 4.33 MW.
Figure 2: Reconfiguration with $K=2$ (line {2, 22} opened and line {24, 28} closed). Bus color shows the voltage level and the percentage of current rating of bottleneck lines are shown. Total DG output is 4.33 MW.
In this reconfiguration, switch $\{2,22\}$ is opened while switch $\{24,28\}$ is closed. This raises the voltages from bus 5 to bus 17, satisfying the lower limit of 0.95 pu (e.g., the voltage at bus 17 is 0.972 pu). If we re-solve (12) with $K=4$, the result is shown in Fig. 3, with total DG output increased to 5.67 MW.
Figure 3: Reconfiguration with $K=4$ (lines {2, 22}, {28, 29} opened, and lines {24, 28}, {17, 32} closed). Bus color shows the voltage level and the percentage of current rating of bottleneck lines are shown. Total DG output is 5.67 MW.
With $K=4$, switch $\{28,29\}$ is opened while switch {17, 32} is closed, in addition to the topology adjustments made in the case of $K=2$. This prevents the two DG sources from sharing the path segment between bus 5 and bus 28, which is the bottleneck for $K=2$. If we further increase $K=6$ in (12), the outcome is to open {10, 11} and close {11, 21}, cutting short the feeder from bus 0 to bus 17. This increases the total DG output to 5.72 MW.
3.2 Computation Time Evaluation
The elapsed runtime to solve the benchmark instances is summarized in Table 2.
Table 2: Runtime [s] to achieve 0.01% optimality gap, or optimality gap [%] after one-hour runtime, for benchmark instances of (12)
name $K=0$ $K=2$ $K=4$ $K=6$ $K=8$
33bw N/A 1.56s 1.69s 2.21s 4.06s
118zh 0.82s 20.18s 28.7s 105.5s 275.74s
136ma 0.14s 4.88s 12s 15.35s 58.16s
REDS 29+1 0.12s 0.62s 0.61s 0.63s 1.38s
REDS 83+11 N/A N/A 0.42s 2.33s 3.32s
REDS 135+1 0.12s 4.46s 19.82s 235.7s 1.99%
REDS 201+3 0.7s 13.43s 41.11s 88.76s 280.96s
REDS 873+7 5s 265s 1337s 3.38% 4.54%
In the table, N/A means that the instance is infeasible. In addition, for REDS 135+1 with $K=8$ and REDS 873+7 with $K=6$ and $K=8$ the table entries show the percentage optimality gap after time limit of one hour. In general, runtime increases as the network size and $K$ increase. However, there are exceptions as REDS 135+1 appears to be difficult for $K=8$ (cf. 136ma with $K=8$ differing only in DG settings), whereas REDS 83+11 is easy (even with $K=26$ the runtime is 11.8s, not shown in Table 2). Despite some irregularities in the runtime pattern, we conclude that the proposed method can reliably obtain guaranteed optimal DG set points and network configuration for moderately-sized instances (e.g., networks up to few hundred buses and $K\leq 6$). However, discretion is needed when solving larger instances. Nevertheless, typically the marginal benefit diminishes rapidly with increasing $K$ (e.g., [15]). Further, despite not being able to solve large instances to optimality, the proposed method can return suboptimal solutions with certified optimality gap. Runtime to achieve 5% optimality gap for the benchmarks is summarized in Table 3.
Table 3: Runtime to achieve 5% optimality gap for benchmark instances of (12)
name $K=0$ $K=2$ $K=4$ $K=6$ $K=8$
33bw N/A 1.14s 1.32s 1.88s 3.92s
118zh 0.31s 14.28s 23.94s 61.93s 119.42s
136ma 0.14s 3.23s 10.87s 14.93s 37.22s
REDS 29+1 0.11s 0.61s 0.61s 0.63s 0.63s
REDS 83+11 N/A N/A 0.27s 1.49s 3.04s
REDS 135+1 0.11s 3.09s 8.80s 5.96s 91.74s
REDS 201+3 0.30s 12.89s 39.41s 84.63s 241.36s
REDS 873+7 0.58s 86.87s 482.4s 1109s 1854s
We conclude that, with reasonable target of optimality gap, the proposed method for DG maximizing ROPF is practical for networks of reasonable size and number of switch status changes. However, for very large number of switch status changes (or a problem to plan for possible line expansion), the proposed method should be augmented with more efficient methodologies.
4 On the Choice of Optimization Solvers and Problem Formulations
This section justifies the choice of our proposed method. We demonstrate that alternative methods for DG maximizing ROPF may not work as intended. These include interior point algorithm, MICP relaxation, linearized DistFlow model, alternative power flow equation formulations and SBB implementations.
4.1 Pitfalls with Local Optimization Solvers
We evaluate MATPOWER’s OPF subroutine runopf.m to solve a special case of (12) with $K=0$. To maximize the total DG output, MATPOWER is set to minimize $\sum_{n}c^{0}_{n}+c^{1}_{n}p_{n}^{G}+c^{2}_{n}(p_{n}^{G})^{2}$ with $c^{0}_{n}=c^{2}_{n}=0$ and $c^{1}_{n}=-1$ for all $n\in\mathcal{N}$. This is not the typical positive quadratic cost function. We run MATPOWER with the default options (e.g., flat start initial guess) and with two optimization solver choices: (a) MIPS interior point method and (b) Matlab’s fmincon active set method. For the feasible instances in Table 2, MATPOWER solves the benchmark OPF instances with relative ease except that MIPS fails to converge for the 201+3 case. However, MATPOWER results are sensitive to problem instance data and algorithmic choices. For example, for benchmark 33bw with line current ratings increased to 600 A for all lines and DG ratings decreased to 8 MW, the OPF instance becomes feasible but MATPOWER results vary greatly depending on the solver choice and initial guess strategy. Table 4 shows the DG maximization results due to the following initial guess strategies [36]: 0 for flat start, 1 for ignoring the system state in the MATPOWER case, 2a and 2b for using the system state in the MATPOWER case (without and with update by first solving (12)) and 3a and 3b for solving power flow and using the resulting state (without and with update by first solving (12)). Table 4 indicates that local optimization methods as in MATPOWER cannot reliably solve the DG maximization problem. On the other hand, the proposed method returns the true optimal DG value of 10.45 MW.
The difficulty exhibited in Table 4 is intrinsic to the DG maximization objective (as in (12)). If we instead minimize $\sum_{n}c^{0}_{n}+c^{1}_{n}p_{n}^{G}+c^{2}_{n}(p_{n}^{G})^{2}$ with $c^{0}_{n}=c^{2}_{n}=0$ and $c^{1}_{n}=1$ (instead of 1-1) for all $n$, MATPOWER can reliably obtain the minimum total DG of 1.93 MW using both MIPS and fmincon for all choices of initial guess strategies.
Table 4: 33bw 600 A, MATPOWER OPF, max DG output
initial guess strategy 0 1 2a
MIPS 9.45 MW 9.45 MW 9.45 MW
fmincon 10.44 MW 10.44 MW fail
initial guess strategy 2b 3a 3b
MIPS 10.44 MW 9.43 MW 10.44 MW
fmincon fail 8.78 MW fail
Therefore, when a ROPF problem with “non-conventional” objective (e.g., DG maximization) is considered, caution must be exercised regarding whether the optimization solver should work as intended.
4.2 Pitfalls with MICP Relaxation
An alternative to (12) is to replace the bilinear equations in (5d) with inequalities, leading to a relaxation [8] as
$\!\!\!\!\underset{\begin{subarray}{c}\nu,\ell,p^{G},q^{G},P,Q,\pi,\alpha\end{subarray}}{\text{maximize}}$ $\quad\sum\limits_{n\in\mathcal{N}}p^{G}_{n}$ (13a)
subject to (2) to (5) except for (5d), (7) to (9) and
$\Big(\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}\nu_{k}\Big)\ell_{n}\geq P_{n}^{2}+Q_{n}^{2},\;\;n\in\mathcal{N}$ (13b)
Unlike (5d), inequalities (13b) allow (13) to be categorized as a MICP problem amenable to efficient solution algorithms. However, a major question regarding (13) is whether it is tight in the sense that (13b) hold as equations at optimality. If (13) is not tight, the control decisions may violate grid constraints (e.g., voltage and current limits) as the DistFlow equations are not satisfied. For (13) with “standard” objective such as total line loss minimization, the tightness of (13) has been empirically and theoretically verified [33, 15, 17, 8]. However, for DG maximization, the issue of tightness has not been explored in depth. In fact, our answer to this question is negative: we will present a three-bus counterexample to show that (13) is not tight and the resulting control will lead to grid constraint violations (this also agrees with the observations in [34]). Consider the OPF problem for the three-bus example in Fig. 4.
Figure 4: 3-bus system showing failure of MICP relaxation. The voltage at bus 0 is 1 pu.
We use Gurobi to solve (12) and (13) with $K=0$. The results are summarized in Table 5, demonstrating that the proposed method is able to obtain the correct optimal DG outputs whereas MICP relaxation fails to do so (the different values of DG outputs are highlighted in red).
Table 5: Per unit OPF results for Fig. 4 by (12) and (13)
Bilinear program in (12)
bus $\nu$ $\ell$ $p^{G}$ $q^{G}$ $P$ $Q$
1 1.1025 25 7.7518 0.39754 4.9991-4.9991 0.092606
2 1.0964 0.2645 0 0 0.50264 0.19736-0.19736
Conic relaxation in (13)
bus $\nu$ $\ell$ $p^{G}$ $q^{G}$ $P$ $Q$
1 1.1025 25 7.9991 0.64489 4.9991-4.9991 0.092606
2 1.0915 25 0 0 0.75 0.049999
With the problem data in Fig. 4 and the variable values in Table 5, we can verify that the solutions of (12) and (13) satisfy all constraints in the respective problems. For example, the reactive power balance equation in (1b) for bus 1 for (12) is $q^{G}_{1}-q^{L}_{1}=0.39754-0.5=-0.10246=(-1)\times 0.092606+(-0.19736)+0.0075\times 25=-Q_{1}+Q_{2}+x_{01}\ell_{1}$. The analogous equation holds for the solution of (13) as well: $q^{G}_{1}-q^{L}_{1}=0.64489-0.5=0.14489=(-1)\times 0.092606+0.049999+0.0075\times 25=-Q_{1}+Q_{2}+x_{01}\ell_{1}$. However, while equation (5d) at bus 2 holds for the solution of (12): $\nu_{1}\ell_{2}=1.1025\times 0.2645=0.2916=0.50264^{2}+{(-0.19736)}^{2}=P_{2}^{2}+Q_{2}^{2}$, (13b) remains an inequality instead of equality for the solution of (13): $\nu_{1}\ell_{2}=1.1025\times 25=27.5625>0.565=0.75^{2}+0.049999^{2}=P_{2}^{2}+Q_{2}^{2}$. Indeed, a posterior load flow analysis based on $(p^{G},q^{G})$ from (13) reveals that the bus voltages are 1, 1.0539 and 1.0511 per unit, respectively. Similarly, the currents for the lines are 522.53 A and 51.235 A. The voltage and current upper limits are both violated by the solution of MICP relaxation. The aforementioned comparison is repeated for all benchmarks in Table 1. MICP relaxation solutions fail to satisfy equality (5d) in all cases, and result in voltage and current violations.
The objective functions in (12) and (13) have a significant impact on the tightness of MICP relaxation. Following [4], we consider the trade-off between DG maximization and line loss minimization by minimizing in (12) and (13) the following weighted sum objective:
$\rho\sum\limits_{n\in\mathcal{N}}\big(\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}r_{kn}\big)\ell_{n}-(1-\rho)\sum\limits_{n\in\mathcal{N}}p^{G}_{n}$ (14)
where $0\leq\rho\leq 1$ is a trade-off parameter. With $\rho$ increasing from 0 to 1, the minimization of (14) transitions from DG maximization to loss minimization. By solving (12) and (13) with (14) and running a posterior load flow analysis based on the resulting DG set points and network configuration, the trade-off can be visualized. Fig. 5 shows the Pareto curves for the 33bw benchmark with universal line rating of 600 A and $K=6$.
Figure 5: Pareto curves of 33bw benchmark (600A line rating) due to the proposed model in (12) and MICP relaxation in (13)
For $\rho>0.5$, the two Pareto curves coincide since MICP relaxation solution satisfies (5d). However, when $\rho\leq 0.5$ MICP relaxation fails as it no longer depicts the true trade-off. This case study raises the awareness that the MICP relaxation approach should be used with caution especially when non-conventional objective function is considered.
4.3 Pitfalls with Linearized DistFlow Approximation
We evaluate a variant of (12) by replacing the DistFlow equations in (5) with its linear approximation (i.e., LinDistFlow equations [3]). The LinDistFlow variant ignores the loss-related terms multiplying $\ell_{n}$ in (5a), (5b) and (5c). Also, equation (5d) is dropped. Substituting the DistFlow equations with the LinDistFlow equations changes (12) to a MILP, which is relatively easy to solve. However, the voltage variables in the LinDistFlow variant overestimate the true voltages (e.g., [30, eq. (2a) and (3)]). Furthermore, since the squared current term $\ell$ is missing it is impossible to impose current upper limit constraints. Indeed, for our benchmark instances severe violation of current upper limit is experienced. To counter this, a surrogate current upper limit constraint can be imposed, similar to (5d), as
$\Big(\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}\nu_{k}\Big)\Big(\sum\limits_{k\in\bar{\mathcal{N}}}\pi_{kn}\overline{\ell}_{kn}\Big)\geq P_{n}^{2}+Q_{n}^{2},\quad n\in\mathcal{N}$ (15)
In (15) the product in the left-hand side can be equivalently replaced by linear terms as illustrated in (6). Below we call LinDistFlow the variant of (12) with LinDistFlow equations and (15). The inclusion of (15) relieves some undervoltage and overcurrent issues, but this is not enough. For instance, for the 33bw benchmark with 600 A line rating and $K=2$, the actual voltages by a posterior load flow analysis and the voltages perceived to be true by the LinDistFlow optimization model (i.e., $(\nu)^{1/2}$) are very different as indicated by Fig. 6.
Figure 6: Benchmark 33bw (600 A line rating): actual voltage profile by a posterior load flow analysis and voltage profile perceived to be true by the LinDistFlow optimization model
Also, current upper limits are violated in multiple lines (worst case violation being 115% of rating) despite the surrogate constraint in (15). This is likely to trigger an undervoltage trip. With reference to (5d), overestimation of voltages will lead to underestimation of currents. Indeed, current upper limit are violated in multiple lines (worst case violation being 115% of rating) despite the surrogate constraint in (15). For LinDistFlow with $K=4$ the result is even more extreme. The load flow analysis by Matpower fails to converge. In contrast, solving (12) for the 33bw 600 A benchmark with $K=2$ and $K=4$, respectively, reliably leads to constraint-abiding control with guaranteed optimality.
4.4 Effect of Power Flow Equation Models
The model of power system physcis (e.g., DistFlow equations versus AC power flow equations) significantly affects the solver’s performance to solve (12). Here we investigate two ROPF models containing power flow equations other than the DistFlow equations (5) adopted in our model in (12):
  • ACOPFC: ROPF model based on standard AC power flow equations by Capitanescu et. al. in [5].
  • ACOPFJ: ROPF model based on modified AC power flow equations by Jabr et. al. in [15].
ACOPFC follows the model in [5]. On the other hand, ACOPFJ is based on the following power system physics model in [15]:
$p^{G}_{n}-p^{L}_{n}=\sum\limits_{k:(n,k)\in\mathcal{L}}P_{nk},$ $n\in\mathcal{N}$ (16a)
$q^{G}_{n}-q^{L}_{n}=\sum\limits_{k:(n,k)\in\mathcal{L}}Q_{nk},$ $n\in\mathcal{N}$ (16b)
$P_{nk}=\sqrt{2}g_{nk}u^{(n,k)}-g_{nk}R_{nk}-b_{nk}T_{nk},$ $(n,k)\in\mathcal{L}$ (16c)
$Q_{nk}=-\sqrt{2}b_{nk}u^{(n,k)}+b_{nk}R_{nk}-g_{nk}T_{nk},$ $(n,k)\in\mathcal{L}$ (16d)
$2u^{(n,k)}u^{(k,n)}\geq R_{nk}^{2}+T_{nk}^{2},$ $(n,k)\in\mathcal{L}$ (16e)
$R_{nk}=R_{kn},$ $(n,k)\in\mathcal{L}$ (16f)
$T_{nk}=-T_{kn},$ $(n,k)\in\mathcal{L}$ (16g)
$0\leq u^{(n,k)}\leq\frac{\overline{\nu}_{n}}{\sqrt{2}}\alpha_{nk},$ $(n,k)\in\mathcal{L}$ (16h)
$0\leq u_{n}-u^{(n,k)}\leq\frac{\overline{\nu}_{n}}{\sqrt{2}}(1-\alpha_{nk}),$ $(n,k)\in\mathcal{L}$ (16i)
where $\mathcal{L}$ is the set of directed lines (normally open and normally closed) such that if $\{n,k\}$ is a line then $(n,k),(k,n)\in\mathcal{L}$, and $g_{nk}+jb_{nk}$ is the series admittance of line $(n,k)$. The decision variables include $P_{nk}$, $Q_{nk}$, $u^{(n,k)}$, $R_{nk}$, $T_{nk}$ and $u_{n}$ for all $(n,k)\in\mathcal{L}$. The ROPF formulation with (16) imposed is a relaxation of (12) because of the inequalities in (16e). However, replacing (16e) with equalities
$2u^{(n,k)}u^{(k,n)}=R_{nk}^{2}+T_{nk}^{2},\quad(n,k)\in\mathcal{L}$ (17)
makes (16) equivalent to the exact AC power flow equations (also equivalent to our DistFlow equations in (5)). The ROPF model with (16) (where the (16e) is replaced by (17)) is the ACOPFJ formulation considered in our case study.
Both ACOPFC and ACOPFJ can be written as MINLP with bilinear equality constraints that is solvable using Gurobi. We test solving these models for the 33bw benchmark using Gurobi with time limit of 300 seconds. Table 6 shows the best DG output (if a feasible solution is found), the percentage optimality gap and the runtime to achieve the gaps. It is evident that for the DG maximizing ROPF, our proposed model in (12) is the only practical choice when compared with ACOPFC and ACOPFJ.
Table 6: Comparing (12), ACOPFC and ACOPFJ, 33bw
33bw $K=2$ $K=4$ $K=6$ $K=8$
(12) 4.34 MW 5.67 MW 5.72 MW 5.72 MW
0%, 1.56s 0%, 1.69s 0%, 2.21s 0%, 4.06s
ACOPFC N/A N/A N/A 4.36 MW
fail, 300s fail, 300s fail, 300s 359%, 300s
ACOPFJ N/A 4.46 MW 4.60 MW 4.60 MW
fail 174%, 300s 202%, 300s 191%, 300s
4.5 Effect of SBB Solver Packages
Different implementations of the SBB algorithm behave very differently when solving (12). Here we evaluate three representative commercial packages:
  • Gurobi 10.0.1: specialized for MINLP allowing only non-convex bilinear equations (and convex constraints).
  • BARON 24.5.8: general SBB based global MINLP solver.
  • Knitro 14.0.0: convex MINLP solver using SBB and interior point algorithms.
The three packages are run with default algorithmic parameter settings. Table 7 shows the comparison result for the benchmark 33bw, showing how the three packages behave differently. The fastest runtime of Gurobi can be explained by the fact that it is specialized to consider only bilinear nonlinearity, whereas BARON is slowest as it aims to handle a much wider class of nonlinearities. Knitro is slower than Gurobi, and it may not return the true optimum as indicated by the MW results for $K=2$ (4.14 MW versus 4.34 MW).
Table 7: Comparing Gurobi, BARON and Knitro, 33bw
33bw $K=2$ $K=4$ $K=6$ $K=8$
Gurobi 4.34 MW 5.67 MW 5.72 MW 5.72 MW
1.56s 1.69s 2.21s 4.06s
BARON 4.34 MW 5.67 MW 5.72 MW 5.72 MW
30.1s 43.0s 68.2s 96.4s
Knitro 4.14 MW 5.67 MW 5.71 MW 5.72 MW
8.64s 8.68s 16.5s 49.5s
The result of a similar comparison case study with the 118zh benchmark is shown in Table 8, also indicating Gurobi’s robustness for the purpose of solving (12).
Table 8: Comparing Gurobi, BARON and Knitro, 118zh
118zh $K=0$ $K=2$ $K=4$ $K=6$
Gurobi 27.47 MW 28.57 MW 29.44 MW 30.17 MW
0.82s 20.2s 28.7s 105.5s
BARON 27.47 MW 28.56 MW 29.27 MW 30.15 MW
2.10s 570s 2964s 2hr (time limit)
Knitro 27.46 MW 28.18 MW 29.35 MW 29.49 MW
0.41s 183s 1034s 1176s
5 Practical Case Study - Real Network
5.1 533-bus Distribution System Description
In this section, the ROPF problem (12) is applied to a real medium voltage distribution system in southern Sweden: the 533-bus network described in [21]. With the expansion of DG, many DSOs wish to assess to what extent DG can be integrated in their systems without violating operational limits. This is referred to as performing a hosting capacity (HC) analysis. In [14], four operational limits that constrain the HC in distribution networks are outlined – voltage, overloading, protection and harmonics. With our proposed method we are imposing voltage and current limits, thereby dealing with the first two. Protection is partly included by requiring radial operation, but no further protection limits (i.e. short-circuit current) are set up. The system is assumed to be three-phase balanced without overtones, thus harmonics mitigation is not considered. Within these limits, the HC can be assessed and enhanced, using e.g. network reconfiguration, and our proposed ROPF method does this jointly.
[…]
Figure 7: Map of the 533-bus distribution system. The width of each line is proportional to its rated current.
The system contains 533 nodes and 577 lines, thus radial operation (3c) requires that 45 lines are always disconnected. Although a disconnected line might still be energized, the existence of an open point renders the entire line topologically switched off. The normal radial topology, used as base case ($K=0$) for subsequent simulations, can be seen in Fig. 8.
The existing wind and solar DG capacity amounts to 34.5 MW, installed at 262 nodes with capacities ranging from 3 kW to 2 MW per node. To avoid curtailment of already existing DG in the system, production from existing DG is embedded in $p^{L}_{n}$ by subtracting hourly generation from hourly load to obtain the net load data for each node. In the year 2022, the total net load in the system varied from 44.6 MW in Dec 16, 08:$0$ to -4.8 MW in Jul 10, 14:$0$. The latter will be referred to as the minimum net load hour.
The voltage constraints in (8) are set to $\sqrt{\underline{\nu_{n}}}=0.95$ pu and $\sqrt{\overline{\nu_{n}}}=1.05$ pu, whereas the current constraints $\overline{\ell_{kn}}$ in (9) are set individually for each line, according to their respective rated currents. All system data was provided by the responsible DSO, Kraftringen, which operates local distribution networks in multiple areas in southern Sweden, including the 533-bus network in question. In the real network, there are often several line segments and conductor types connecting two substations, while in the model these are aggregated to single lines with equivalent rated currents and impedances. A full description of this aggregation procedure, and the model construction overall, can be found in [21]. Once the aggregation has been made, the single lines have rated phase currents ranging from 93-1200 A, phase resistances from 0.003-1.86 $\Omega$ and phase reactances from 0.001-0.73 $\Omega$.
Furthermore, the real network is a three-phase system, whereas the optimization algorithm is designed for single-phase. Since the system is assumed to be three-phase balanced, this was managed by dividing all loads and generation capacities by a factor three, effectively treating each phase as a separate system. After simulations were run, the actual power flows were obtained by multiplying all power values post-optimization with a factor three. The final Matpower-model of the 533-bus system can be found at github.com/MATPOWER/matpower/blob/master/data/case533mt_lo.m with net load values from the minimum net load hour.
5.2 Reconfiguration with 533-bus Example
The ROPF problem in (12) is solved for the 533-bus system through two separate case studies. The first study involves the construction of a multiple node generation scenario, representing a potential future expansion of wind power. The second study is a spatially comprehensive DG expansion site analysis where additional DG capacity is added only to a single node in the system one at a time but iterating through the entire system, to see how high the HC is in each specific location and how much it can be increased through network reconfiguration.
In both the wind scenario and the DG expansion site analysis, the problem is solved with the minimum net load hour of 2022 as load input. The minimum net load hour is considered to be a good worst-case scenario, since bus voltages throughout the network are high during this hour and overvoltage is frequently the HC limiting factor.
Additionally, in both the wind scenario and the DG expansion site analysis, results are given for $K=0$, 22 and 44 respectively. Naturally, the method can be applied also with higher $K$-values. However, this usually renders quickly diminishing marginal returns as shown in [15], [21] and Sec. 5.2.1. In other words, the highest increases in HC are generally gained from the first switch events. Moreover, the optimization time rapidly increases as the search space grows exponentially with $K$. Finally, DSOs are in practice seldom interested in making too drastic topology changes in their networks based on specific load scenarios. The low $K$-scenarios reveal small topology changes from the normal configuration with potentially high rewards, which is attractive for grid planners.
5.2.1 Wind scenario
In the wind scenario, DG units of 2 MW are added to 21 rural nodes in clusters of up to four units. The placement of these clusters can be seen in Fig. 8. The clusters are added without grid expansion or reinforcement, to determine the HC of the existing grid, before and after reconfiguration. Note that the 21 wind units are the only $p^{G}_{n}$ in the optimization, since existing DG is embedded in $p^{L}_{n}$.
Figure 8: Setup of the wind scenario in the normal radial topology. Green circles display the 21 nodes where additional DG units of 2 MW have been placed.
Solving (12) for this generation setup, $\sum_{n}p^{G}_{n}=17.6$, $19.2$ and $19.3$ MW for $K=0$, 22 and 44 respectively. As an initial observation, the added DG capacity of $21\times 2=42$ MW is significantly curtailed both with and without reconfiguration. With $K=2$, the HC increases by 1.6 MW (or 9.1%) compared to the normal topology. This increase is due to a pair of switch events that divides a cluster of four units in the southeast on two separate feeder lines, in contrast to the normal topology where they share a common feeder that turns out to be a bottleneck (see Fig. 9). If four switch events are allowed ($K=4$), the algorithm returns the same first pair of switch events as with $K=2$, but also another pair of switch events closer to the distribution station LF. This has a more limited effect, increasing HC by only 0.1 MW from $K=2$.
Figure 9: Left: Wind scenario in the normal radial topology ($K=0$), zoomed in at the southeastern part of the network. Right: Wind scenario after reconfiguration ($K=2$), the black arrows indicate where switch events occurred.
5.2.2 DG expansion site analysis
In the DG expansion site analysis, additional DG capacity is added to a single node one at a time. This added capacity is large enough for the apparent power rating of the unit to not be an active constraint in (7b), allowing us to identify the HC of the grid. After adding DG to a single node, (12) is solved with $K=0$, 22 and 44. This procedure is repeated for all 532 non-slack nodes in the system, sometimes referred to as the iterative or detailed method [31]. No single optimization took more than 10 mins, which is considered reasonable for a site analysis.
In Fig. 10, the results from the $532\times 3$ optimizations in the DG expansion site analysis can be seen. The $x,y$-value indicates the location of a node in the network and the bar height is the HC at that node in MW. The blue part of each bar is the HC in the normal radial topology, whereas the red and yellow part is the additional HC including reconfiguration with $K=2$ and 44 respectively.
Figure 10: 3D bar chart displaying the HC at each node. The blue, red and yellow parts are with $K=0$, 22 and 44 respectively.
Tall bars in Fig. 10 indicate a high HC, typically found at distribution substations or large industry nodes. On average, the HC is also higher at urban nodes than at rural. A large share of red or yellow in a bar indicates that the HC at that node can be improved substantially by reconfiguration. A large share of the urban nodes, but also some rural, exhibit this substantial increase in HC from reconfiguration. In Fig. 11, these relative HC improvements from network reconfiguration are presented explicitly. The figure displays how many of the nodes, $x\%$, that exhibit a relative HC increase of $y\%$ or higher from reconfiguration.
Figure 11: Share of nodes, $x\%$, that exhibit a HC increase of $y\%$ or higher from reconfiguration. Red depict the increase from $K=0$ to $K=2$, yellow from $K=0$ to $K=4$.
Fig. 11 shows that the potential to enhance the HC through network reconfiguration varies considerably between nodes. The location of a node, and the network topology in its vicinity, has a significant effect on the benefit that can be achieved from reconfiguration. The greatest increase from $K=0$ to $K=4$ is as high as 212% for a single node. Meanwhile, the median increase from $K=0$ to $K=4$ is only about 11%.
6 Conclusion and Future Work
DG maximization is more difficult than loss reduction for ROPF as the former objective renders standard techniques such as local optimization, MICP relaxation and linearization approximation ineffective, thus raising caution against their over-generalization and highlighting the importance of decision models featuring exact AC power flow equations (or equivalents). With a proper choice of problem formulation and optimization software package (e.g., DistFlow equations and Gurobi SBB), DG maximizing ROPF can indeed be handled reasonably well for moderately sized networks in a time scale relevant even for control center applications.
In this paper, the three-phase balanced problem has been considered. This is a natural starting point and a reasonable generalization as many distribution networks across the world are in balanced or close to balanced operation, the networks using three-phase feeders instead of single-phase laterals [22],[25]. For future work however, attention should be paid also to the unbalanced case, which is a reality in many networks and an important source of complexity. The ability of DG inverters and other sources of voltage control to compensate unbalances between phases could then be exploited further. Indeed, there exist variants of the DistFlow equations for three-phase unbalanced and meshed networks. Incorporating these equations in the ROPF model and identifying the proper solution approach are worth further investigation.